A sample of iodine has a mass of 36.5 grams. How many I atoms are there in the sample? How many I2 molecules are there in the sample?  molecules

A sample of chlorine has a mass of 37.0 grams. How many  atoms are there in the sample? How many Cl2 molecules are there in the sample?
Answer #1

1)

Molar mass of I2 = 253.8 g/mol

mass(I2)= 36.5 g

use:

number of mol of I2,

n = mass of I2/molar mass of I2

=(36.5 g)/(2.538*10^2 g/mol)

= 0.1438 mol

use:

number of molecules = number of mol * Avogadro’s number

number of molecules = 0.1438 * 6.022*10^23 molecules

number of molecules = 8.66*10^22 molecules

Count of each element in compound is:

{I=2}

Total number of atoms in 1 molecule = 2

use:

number of atoms = number of molecules * Total number of atoms in 1 molecule

number of atoms = 8.66*10^22 * 2

number of atoms = 1.732*10^23

Answer: 1.73*10^23

2)

Molar mass of Cl2 = 70.9 g/mol

mass(Cl2)= 37.0 g

use:

number of mol of Cl2,

n = mass of Cl2/molar mass of Cl2

=(37 g)/(70.9 g/mol)

= 0.5219 mol

use:

number of molecules = number of mol * Avogadro’s number

number of molecules = 0.5219 * 6.022*10^23 molecules

number of molecules = 3.143*10^23 molecules

Count of each element in compound is:

{Cl=2}

Total number of atoms in 1 molecule = 2

use:

number of atoms = number of molecules * Total number of atoms in 1 molecule

number of atoms = 3.143*10^23 * 2

number of atoms = 6.285*10^23

Answer: 6.28*10^23